# Lijuan Wang, Yao Qian, Jinyu Li Microsoft

### v1

@@ Abstract
Index Terms:
Speech language models, conversational AI
@@ II-A Fair comparison of decoding paradigms
The comparison of different patterns are shown in Table I. Although decoding paradigms form the foundation of speech language models, comparing these strategies in a controlled setting remains challenging. Key difficulties include variations in training data, training recipes, base language model configurations, speech tokenization methods, and the number of trainable parameters. To ensure a fair and meaningful comparison, we standardize most of critical components to compare the jointly speech-text modeling patterns within one language model, include the interleaved and the parallel patterns: using the same base language model, (Phi-4 MM [17]), the same supervised fine-tuning (SFT) data, the same number of trainable parameters, and the same speech tokenization method (S3Tokenizer [18]). In the Thinker-Talker decoding paradigm, the Talker is decoupled from the Thinker and functions similarly to a pseudo TTS module—taking text tokens and hidden states from the language model to generate speech tokens. This separation prevents speech token prediction from interfering with the Thinker’s behavior, thereby preserving the original language model’s text capabilities. Since our primary focus is on jointly trained decoding paradigms within one LM, where the main language model both predicts and consumes text and speech tokens, we include the Thinker-Talker approach only as a reference for comparison [14, 15]. We use the officially released checkpoints of LLaMA-Omni-2 for comparison, as its training setup closely aligns with ours—train on only a few thousand speech QA data, unlike other models trained on millions of diverse datasets. Additionally, LLaMA-Omni-2 was released recently (May 5, 2025), making it a timely and relevant baseline. We argue that it still serves as a reasonable baseline for comparing with other decoding patterns.
@@ II-B2 Proposed early-stop interleaved decoding paradigm
[…]
@@
II-B3 System overview of Phi-Omni-ESI
Figure 2 presents an overview of our Phi4-Omni-ESI, a speech-in, speech-out question answering system, built based on a multimodal open-source language model (Phi4-MM [17]). The input text is consist of system prompts or user instructions. The audio question is first processed by an audio encoder—initialized from a pretrained ASR encoder and further fine-tuned on ASR data—to produce speech embeddings. These embeddings, together with the text input, are fed into the main speech language model. We use the S3Tokenizer [18] to extract the speech tokens. The model performs decoding using our proposed early-stop decoding pattern, generating both text and speech tokens. The generated text can be directly presented to the user as the textual answer, while the speech tokens are passed to a streaming audio token vocoder [18], to synthesize the speech response.
@@ III-2 Model Configuration
We use the open-source Phi4-MM [^6] https://huggingface.co/microsoft/Phi-4-multimodal-instruct speech encoder as our audio encoder. For the language model, we adopt the Phi4-MM series, including the 3.8B and 7B variants. The speech tokenizer [^7] https://github.com/xingchensong/S3Tokenizer, audio token vocoder (including both the flow matching model and neural vocoder) are directly adopted from CosyVoice 2 [18] [^8] https://github.com/FunAudioLLM/CosyVoice without modification.
@@ III-3 Training details
In line with the configuration described in the Phi-4-Mini paper [17], we incorporate a Low-Rank Adaptation (LoRA) [19] module with a rank of 320. Our experiments are conducted using two base language models, with 3.8 billion and 7 billion parameters, respectively. When equipped with LoRA, the total number of trainable parameters becomes 460 million and 707 million for the 3.8B and 7B models, respectively. Rather than updating the full model, we fine-tune only the LoRA layers.
@@ IV-A Main results
Comparing our methods with models trained on millions of hours data: (1). GLM-4-Voice is one of the earliest models to adopt the interleaved speech-text decoding strategy. Our proposed method, Phi-Omni-ESI, extends and improves this paradigm. On all benchmarks, Phi-Omni-ESI 7B consistently outperforms GLM-4-Voice, particularly in speech output accuracy and speech-to-text alignment (relative percentage). Notably, Phi-Omni-ESI 3.8B surpasses GLM-4-Voice on the LLaMA Questions (and is on par on the Web Questions), despite using a significantly smaller base model (3.8B vs. 9B). (2). Compared to Moshi, Phi-Omni-ESI 7B also achieves stronger and more stable performance across tasks. Even the smaller 3.8B variant (referred to as C1) outperforms Moshi in all metrics. (3). Our model C2 exceeds Minmo in both speech output accuracy and relative percentage, demonstrating more effective alignment between generated speech and text. Note that our base models, C0.1 and C0.2, do not outperform Minmo on text-based QA. Minmo is with a better base model with outstanding intelligence than ours. We believe that replacing the base model could lead to significant performance improvements for our approach.
Unlike models such as GLM-4-Voice, Moshi, and Minmo, which are trained on millions of diverse examples, LLaMA-Omni-2 and Phi-Omni-ESI are both fine-tuned using only a few thousand hours of QA samples. This makes them more comparable. When comparing Phi-Omni-ESI 3.8B and 7B to the LLaMA-Omni-2 3B, 7B, and 14B counterparts: (1). Phi-Omni-ESI shows clear advantages in speech-out accuracy, Rel %, and WER. Take LLaMA Questions as an example, Phi-Omni-ESI 3.8B achieves 64.67% speech-out accuracy, surpassing LLaMA-Omni2-3B (58.33%) and even LLaMA-Omni2-7B (64.0%). Its relative score of 0.95 further indicates a high degree of alignment between text and speech outputs. Moving to the 7B variant, Phi-Omni-ESI 7B pushes speech-out accuracy to 65%, outperforming all LLaMA-Omni2 variants, including the much larger 14B model (63.33%). (2). When comparing models of similar size, Phi-Omni-ESI 3.8B outperforms LLaMA-Omni2-3B across all tasks. For example, On LLaMA-QA, it achieves 64.67% speech-out accuracy versus 58.33% for LLaMA-Omni2-3B, and shows stronger relative alignment (0.95 vs. 0.87). At the 7B scale, Phi-Omni-ESI 7B also consistently outperforms LLaMA-Omni2-7B for a large margin for all metrics on Web Questions and Trivia-QA. on LLaMA questions, Phi-Omni-ESI 7B continues to outperform LLaMA-Omni2-7B, achieving 65% vs. 64% for the speech-out accuracy, and showing higher relative scores.
@@ IV-C ESI pattern vs Interleaved pattern
As shown in Table V, we compare interleaved decoding (I) and early-stop interleaved decoding (ESI) strategies using the same base models—3.8B and 7B—on the LLaMA Questions and Web Questions benchmarks. In this table, “V” denotes training only on the VoiceAssistant QA dataset, while “All” includes additional data from Trivia-QA and NaturalQuestions. As results for Trivia-QA are consistent, we omit them here for limited spare space.

### v3

@@ II-A Fair comparison of decoding paradigms
The comparison of different patterns are shown in Table I. Although decoding paradigms form the foundation of speech language models, comparing these strategies in a controlled setting remains challenging. Key difficulties include variations in training data, training recipes, base language model configurations, speech tokenization methods, and the number of trainable parameters. To ensure a fair and meaningful comparison, we standardize most of critical components to compare the jointly speech-text modeling patterns within one language model, include the interleaved and the parallel patterns: using the same base language model, (Phi4- MM [13]), the same supervised fine-tuning (SFT) data, the same number of trainable parameters, and the same speech tokenization method (S3Tokenizer [7]). In the Thinker-Talker decoding paradigm, the Talker is decoupled from the Thinker and functions similarly to a pseudo TTS module—taking text tokens and hidden states from the language model to generate speech tokens. This separation prevents speech token prediction from interfering with the Thinker’s behavior, thereby preserving the original language model’s text capabilities. Since our primary focus is on jointly trained decoding paradigms within one LM, where the main language model both predicts and consumes text and speech tokens, we include the Thinker-Talker approach only as a reference for comparison [3, 8]. We use the officially released checkpoints of LLaMA-Omni-2 for comparison, as its training setup closely aligns with ours—train on only a few thousand speech QA data, unlike other models trained on millions of diverse datasets. Additionally, LLaMA-Omni-2 was released recently, making it a timely and relevant baseline. We argue that it still serves as a reasonable baseline for comparing with other decoding patterns.
@@ II-B2 Proposed early-stop interleaved decoding paradigm
To address the inefficiency caused by long interleaved sequences, as shown in Figure 2, we introduce an early-stop interleaved decoding mechanism that enables the model to skip unnecessary text padding tokens. We construct interleaved sequences using a fixed text-to-speech token ratio of 5:10 before the text tokens are exhausted. Once the model finishes generating all meaningful text tokens and predicts the end-of-sentence (EOS) token for text, then a special token ($\text{S}>$) is inserted to indicate that all remaining tokens in the sequence will be speech tokens [^2] In our preliminary experiments, we observed that without the special token $\text{S}>$, the language model struggled to initiate audio token generation. This token acts as a marker to signal the start of the remaining audio sequence, prompting the model to switch into a mode where all subsequent tokens are treated as audio tokens.. This design enables the removal of redundant text padding tokens that would otherwise be inserted to maintain a fixed interleaving ratio. For example, with a text-to-speech token ratio of 5:10, eliminating padding reduces the total sequence length to approximately 75% of its original length, resulting in lower computational overhead and enhancing both inference speed and training efficiency.
@@
II-B3 System overview of Phi-4MM with ESI
Figure 2 presents an overview of our Phi4-MM with ESI, a speech-in, speech-out question answering system, built based on a multimodal open-source language model (Phi4-MM [13]). The input text is consist of system prompts or user instructions. The audio question is first processed by an audio encoder—initialized from a pretrained ASR encoder and further fine-tuned on ASR data—to produce speech embeddings. These embeddings, together with the text input, are fed into the main speech language model. We use the S3Tokenizer [7] to extract the speech tokens. The model performs decoding using our proposed early-stop decoding pattern, generating both text and speech tokens. The generated text can be directly presented to the user as the textual answer, while the speech tokens are passed to a streaming audio token vocoder [7], to synthesize the speech response.
@@ III-2 Model Configuration
We use the open-source Phi4-MM [^6] https://huggingface.co/microsoft/Phi-4-multimodal-instruct speech encoder as our audio encoder. For the language model, For the language model, we adopt Phi4-MM 3.8B as the base model and further scale it up to 7B. The speech tokenizer [^7] https://github.com/xingchensong/S3Tokenizer, audio token vocoder (including both the flow matching model and neural vocoder) are directly adopted from CosyVoice 2 [7] [^8] https://github.com/FunAudioLLM/CosyVoice without modification.
@@ III-3 Training details
In line with the configuration described in the Phi4-MM paper [13], we incorporate a Low-Rank Adaptation (LoRA) [10] module with a rank of 320. Our experiments are conducted using two base language models, with 3.8 billion and 7 billion parameters, respectively. When equipped with LoRA, the total number of trainable parameters becomes 460 million and 707 million for the 3.8B and 7B models, respectively. Rather than updating the full model, we fine-tune only the LoRA layers.
@@ IV-A Main results
Comparing our methods with models trained on millions of hours data: (1). GLM-4-Voice is one of the earliest models to adopt the interleaved speech-text decoding strategy. Our proposed method, Phi4-MM with ESI, extends and improves this paradigm. On all benchmarks, Ours 7B-ESI consistently outperforms GLM-4-Voice, particularly in speech output accuracy and speech-to-text alignment (relative percentage). Notably, Ours 3.8B-ESI surpasses GLM-4-Voice on the LLaMA Questions (and is on par on the Web Questions), despite using a significantly smaller base model (3.8B vs. 9B). (2). Compared to Moshi, Ours 7B-ESI also achieves stronger and more stable performance across tasks. Even the smaller 3.8B variant (referred to as C1) outperforms Moshi in all metrics. (3). Our model C2 exceeds Minmo in both speech output accuracy and relative percentage, demonstrating more effective alignment between generated speech and text. Note that our base models, C0.1 and C0.2, do not outperform Minmo on text-based QA. Minmo is with a better base model with outstanding intelligence than ours. We believe that replacing the base model could lead to significant performance improvements for our approach.
Unlike models such as GLM-4-Voice, Moshi, and Minmo, which are trained on millions of diverse examples, LLaMA-Omni-2 and Phi4-MM with ESI are both fine-tuned using only a few thousand hours of QA samples. This makes them more comparable. When comparing Ours-3.8B ESI and 7B-ESI to the LLaMA-Omni-2 3B, 7B, and 14B counterparts: (1). Phi4-MM with ESI shows clear advantages in speech-out accuracy, Rel %, and WER. Take LLaMA Questions as an example, Ours 3.8B-ESI achieves 64.67% speech-out accuracy, surpassing LLaMA-Omni2-3B (58.33%) and even LLaMA-Omni2-7B (64.0%). Its relative score of 0.95 further indicates a high degree of alignment between text and speech outputs. Moving to the 7B variant, Ours 7B-ESI pushes speech-out accuracy to 65%, outperforming all LLaMA-Omni2 variants, including the much larger 14B model (63.33%). (2). When comparing models of similar size, Ours 3.8B-ESI outperforms LLaMA-Omni2-3B across all tasks. For example, On LLaMA-QA, it achieves 64.67% speech-out accuracy versus 58.33% for LLaMA-Omni2-3B, and shows stronger relative alignment (0.95 vs. 0.87). At the 7B scale, Ours 7B-ESI also consistently outperforms LLaMA-Omni2-7B for a large margin for all metrics on Web Questions and Trivia-QA. on LLaMA questions, Ours 7B-ESI continues to outperform LLaMA-Omni2-7B, achieving 65% vs. 64% for the speech-out accuracy, and showing higher relative scores.
@@ IV-C ESI pattern vs Interleaved pattern
As shown in Table V, we compare interleaved decoding (I) and early-stop interleaved decoding (ESI) strategies using the same base models 3.8B and 7B on the LLaMA Questions and Web Questions benchmarks. In this table, “V” denotes training only on the VoiceAssistant QA dataset, while “All” includes additional data from Trivia-QA and NaturalQuestions. As results for Trivia-QA are consistent, we omit them here for limited spare space.